(y ^ 2 * sin x) * d * x + (1 / x - y / x) * d * y = 0
1
Expert's answer
2021-09-28T15:51:28-0400

"y^2\sin xdx+\frac{1-y}{x}dy=0\\
\frac{1-y}{y^2}dy=-x\sin xdx\\
\text{Integrate both sides}\\
\int \frac{1}{y^2}-\frac{1}{y}dy=-\int x\sin x dx\\
-\frac{1}{y}-\ln y=-[-x\cos x-\int -\cos x dx]\\
-\frac{1}{y}-\ln y=x\cos x-\sin x +C\\
e^{\frac{1}{y}+\ln y}=e^{-[x\cos x-\sin x +C]}\\
ye^{\frac{1}{y}}=e^{-[x\cos x-\sin x +C]}"